A 300N wagon is pulled up a 30° incline at constant speed.


a. How large a force parallel to the incline is needed if friction effects total of 10N?

b. What will be the force exerted by the incline to support the wagon?

c. What is the coefficient of friction?


1
Expert's answer
2021-01-22T09:57:17-0500

a)


"F=300\sin{30}+10=160\ N"

b)


"N=300\cos{30}=260\ N"

c)


"\mu=\frac{10}{260}=0.04"